### 22 Feb 2024 v1

# A Decision-Language Model (DLM) for Dynamic Restless Multi-Armed Bandit Tasks in Public Health

Abstract
Efforts to reduce maternal mortality rate, a key UN Sustainable Development target (SDG Target 3.1), rely largely on preventative care programs to spread critical health information to high-risk populations. These programs face two important challenges: efficiently allocating limited health resources to large beneficiary populations, and adapting to evolving policy priorities. While prior works in restless multi-armed bandit (RMAB) demonstrated success in public health allocation tasks, they lack flexibility to adapt to evolving policy priorities. Concurrently, Large Language Models (LLMs) have emerged as adept, automated planners in various domains, including robotic control and navigation. In this paper, we propose DLM: a Decision Language Model for RMABs. To enable dynamic fine-tuning of RMAB policies for challenging public health settings using human-language commands, we propose using LLMs as automated planners to (1) interpret human policy preference prompts, (2) propose code reward functions for a multi-agent RL environment for RMABs, and (3) iterate on the generated reward using feedback from RMAB simulations to effectively adapt policy outcomes. In collaboration with ARMMAN, an India-based public health organization promoting preventative care for pregnant mothers, we conduct a simulation study, showing DLM can dynamically shape policy outcomes using only human language commands as input.
1 Introduction
Maternal and child health is a critical public health priority worldwide. Recent estimates indicate that the global maternal mortality ratio is more than double the UN Sustainable Development Goal target (SDG Target 3.1) of below 70 deaths per 100k live births United Nations and Development (2015). Maternal mortality (MM) is of particular concern for low and middle-income countries, which account for 99% of global maternal deaths Kurjak et al. (2023). Further, a large portion of MMs stem from preventable causes Organization and others (2015); as such, programs promoting preventative care awareness, for example through mobile health platforms, have shown to considerably improve access to essential care information Choudhury et al. (2021).
However, two key problems remain for these preventative care programs, often operated by non-profits supporting low-resource communities HelpMum (2023); ARMMAN (2019). First, these programs typically operate with insufficient financial and human resources, highlighting the need for effective resource allocation strategies to maximize health outcomes Kruk et al. (2018); Baltussen and Niessen (2006). Second, these programs must frequently adapt to changing population needs and intervention priorities Organization and others (2015). Adaptability is especially important for prioritizing care for known subpopulations with higher risk Gräler et al. (2022), and for adjusting strategies using community or expert-driven insights Deardorff et al. (2018).
Restless multi-armed bandits (RMAB) have been shown to be successful in modelling resource allocation in public health Nishtala et al. (2023); Wang et al. (2023). In the classical RMAB formulation, a central planner chooses arms to allocate resources and observes state-dependent reward from arms. Despite the strength of RMABs in planning resource allocation, a significant challenge persists in developing models capable of adapting to changing policy objectives. For example, public health experts may wish to use RMAB models to direct resources towards specific underprivileged subpopulations. However, existing works in RMABs focus on fixed policy objectives and require significant human tuning and re-training to meet new desired health outcome.
Recently, large language models (LLMs) have emerged as powerful planners in a wide variety of tasks including navigation Shah et al. (2023), spatio-temporal reasoning Vemprala et al. (2023), and interactive decision-making Li et al. (2022). Recent works have also shown that LLMs can be used directly as code generators to propose reward functions for reinforcement learning (RL) to complete complex robotic manipulation tasks Ma et al. (2024). However, the ability of LLMs as planners for resource allocation in real-world public health settings remains unstudied. This planning ability is especially useful for non-profits that have evolving priorities but do not have the capacity to frequently retrain models.
We propose a Decision Language Model for RMABs, combining strengths of LLMs in semantic comprehension, code generation, and in-context learning, with the efficiency of RMAB modeling in resource allocation for public health. To enable rapid policy adjustment using human language commands, we propose: 1) using LLMs to disambiguate human-expressed policy preferences, 2) using LLMs to directly propose reward functions as code to reach desired RMAB objectives, and 3) a mechanism for iteratively refining LLM-generated reward functions without any ground truth feedback. Using simulation results of LLM-proposed reward functions, we continually improve the alignment of reward functions towards specified goals.
We partner with the India-based non-profit ARMMAN, a public health organization that spreads preventative care awareness to pregnant mothers through an automated phone-based messaging system. To increase retention, ARMMAN employs health workers to provide service calls to beneficiaries. Due to limited support staff, this assignment is similar to a limited resource allocation problem. Prior works have used RMABs to model the user retention problem in the ARMMAN setting Killian et al. (2023); Wang et al. (2023), but they fail to address evolving priorities of ARMMAN.
In summary, our key contributions are as follows:
  • To the best of our knowledge, we are the first to propose using LLMs to adapt to changing resource allocation objectives in public health through reward design.
  • We introduce a reward proposal loop that enhances LLM-generated reward functions using feedback from restless multi-armed bandit (RMAB) simulations, enabling LLMs to iteratively refine reward design to achieve specific, human-specified policy outcomes.
  • To assess the feasibility of deploying such a system in the real world, we evaluate our algorithms’ performance in a real-world task of resource allocation for maternal and child care, and demonstrate near human-level reward-design performance. Our algorithm enables dynamically adjusting policies to match human-specified outcomes, using only language prompts as input.
2 Related Work
RMABs The restless multi-armed bandit problem introduced by Whittle Whittle (1988), is classically solved through the Whittle index policy Weber and Weiss (1990); Glazebrook et al. (2006). Subsequent works have generalized to multi-action RMABs Glazebrook et al. (2011); Killian et al. (2022). RMABs gained prominence in public health domains and are deployed to disseminate preventative healthcare information, monitor health program adherence, and model disease spread Tambe (2022); Mate et al. (2022). However, existing works focus on fixed reward functions and fail to consider that public health planners often have evolving priorities Organization and others (2015).
Reward Design: Designing reward functions that effectively condense long-term agent goals into immediate behavioral signals is a central problem in RL Singh et al. (2009). Manual designs are prone to task misspecification and overfitting Booth et al. (2023). Ho and Ermon Ho and Ermon (2016) and Hussein et al., Hussein et al. (2017) reshape reward through expert behavior observation, but they ignored that expert examples may not be available in resource and data-limited public health domains. Using LLMs as dense reward signals have been investigated Kwon et al. (2022); Du et al. (2023). Ma et al. Ma et al. (2024) use LLMs to output reward function code from language-specified goals. However, above works failed to address multi-agent settings, where a central planner could prioritize certain groups through reward design.
3 Background
We consider an RMAB problem with $N$ arms, representing public health program beneficiaries. Each arm $n\in[N]$ follows a Markov decision process $(\mathcal{S}_{n},\mathcal{A}_{n},\mathcal{C}_{n},T_{n},R_{n},\beta_{n})$ with state space $\mathcal{S}_{n}$, action space $\mathcal{A}_{n}$, action costs $\mathcal{C}_{n}:\mathcal{A}_{n}\rightarrow\mathbb{R}$, and unknown transition probabilities $T_{n}:\mathcal{S}_{n}\times\mathcal{A}_{n}\times\mathcal{S}_{n}\rightarrow[0,1]$. We define reward function $R_{n}:\mathcal{S}_{n}\rightarrow\mathbb{R}$ and discount factor $\beta\in[0,1)$. When $\mathcal{S}_{n}$, $\mathcal{A}_{n}$, and $\mathcal{C}_{n}$ are the same for all arms $n\in[N]$, as in the public health setting we focus on, we omit the subscript $n$. Note our algorithms also applie to more general settings. A planner learns a policy $\pi$ that maps states $\bm{s}\in\mathbb{R}^{N}$ to actions in the set $\mathcal{A}$, given the constraint that the sum cost of actions does not exceed budget $B$. The objective is to learn a policy that maximizes the Bellman equation:
$J(\bm{s})=\max_{\pi\in\Pi}\left\{\sum_{n=1}^{N}R\left(\bm{s}_{n}\right)+\beta\,\mathbb{E}\left[J\left(\bm{s}^{\prime}\right)\mid\bm{s},\pi\right]\right\}\,,$ (1)
[…]
for cost of action $c_{\pi(s_{n})}\in\mathcal{C}$. We take the Lagrangian relaxation to decouple the value functions across arms:
$J\left(s,\lambda^{\star}\right)$ $=\min_{\lambda\geq 0}\left(\frac{\lambda B}{1-\beta}+\sum_{n=1}^{N}\max_{a_{n}\in\mathcal{A}}\left\{Q_{n}\left(\bm{s}_{n},a_{n},\lambda\right)\right\}\right),$ (3)
s.t. $Q_{n}\left(\bm{s}_{n},a_{n},\lambda\right)=R\left(\bm{s}_{n}\right)-\lambda c_{a_{n}}$
$+\beta\,\mathbb{E}\left[Q_{n}\left(\bm{s}_{n}^{\prime},a_{n},\lambda\right)\mid\pi(\lambda)\right].$
for Q-function $Q$, arm $n$ action $a_{n}$, transitioned state from $s_{n}$ under action $a_{n}$ as $s_{n}^{\prime}$, and $\pi(\lambda)$ an optimal policy under $\lambda$. Shinn et al. (2024); Ma et al. (2024). We build on deep RL approaches for RMABs, which we discuss in Appendix C.
Figure 1: Overview of the proposed DLM loop. We begin by providing three key context descriptions to the LLM. First, we provide a language command; we include the specific commands used in Appendix Table 1. We also provide arm feature context, which contains a list of relevant per-arm features available to use in the proposed reward, as well as syntax cues and the relevant subset of the codebase so that the proposed LLM reward functions may be used directly in code. From this context, the LLM then proposes candidate reward functions which are used to train optimal policies under proposed reward. These trained policies are then used to generate policy outcome comparisons, which show the distribution of rewards across key feature groups in simulation. Finally, we query an LLM to perform self-reflection Shinn et al. (2024); Ma et al. (2024) by choosing the best candidate reward aligning with the original language command; the selected best rewards are used as context for future reward generation iterations.
4 Decision-Language Model for RMABs
Below, we provide an overview of the proposed DLM reward generation technique. An overview of the proposed technique is provided in Fig. 1. We first describe the problem setting of reward generation, and motivate the use of LLMs for reward generation from a theoretical perspective. We then discuss the context is provided to the LLM for initial reward candidate generation. Then, we discuss the independent simulation stage which identifies an optimal policy $\pi^{R*}$ under a proposed reward $R$, and the subsequent policy outcome comparison stage that simulates the respective policy reward outcomes. To be clear, $\pi^{R*}$ references the optimal policy induced by a given reward function $R$. This does not refer to an optimal reward $*R*$. Finally, we discuss the LLM reflection stage which considers the policy outcome comparison and identifies successful candidates to guide future in-context learning for LLM reward generation.
4.1 Problem Setting: Reward Generation for RMABs
We first define the goal of language-conditioned reward design in the restless multi-armed bandit setting. We consider a human-specified language policy goal $\ell$ with a corresponding base reward function that maps policies $\pi\in\Pi$ to real, scalar values: $F:\Pi\rightarrow\mathbb{R}$. This base reward function establishes a scalar “ground truth” evaluation that corresponds directly to the human-specified command. Evaluation through the base reward is available through real or simulated queries of the proposed policy. Then, our objective is to optimize $R$:
$\max_{R\in\mathcal{R}}F(\pi^{R*}),\text{where: }$:
$\pi^{R*}=$ $\mathop{\mathrm{argmax}}_{\pi\in\Pi}\left\{\sum_{n=1}^{N}R\left(\bm{s}_{n},\bm{z}\right)+\beta\,\mathbb{E}\left[J\left(\bm{s}^{\prime}\right)\mid\bm{s},\pi\right]\right\}$
[…]
In summary, our key objective is, given a human-specified language prompt $\ell$, to generate the reward function $R$ that maximizes base function $F$ through optimal policy $\pi^{R*}$ for reward $R$. We propose using LLMs to automatically propose this reward function $R$, and automatically refine proposed rewards through an iterative self-reflection procedure.
4.2 Sample Inefficiency of Fitness Function
In order to evaluate each task description that we provide, we design a ‘ground truth fitness function’. This is the reward function (assigning a reward to each state and action) against which the learned policies are evaluated. However, this fitness function is unknown to the LLM. The ‘base’ policy in our comparison is obtained by training an RL policy to maximize the discounted reward obtained by this ground truth fitness function and serves as a baseline to our experiments. We want to note that, despite being the ‘true reward function’, this might not be the best reward function to learn the optimal policy in a learning theoretic sense. This is usually the case when rewards are sparse (i.e, hard to obtain) from simple policies and the sparsity of the reward signal makes learning hard. Thus, reward shaping/ reward engineering has been extensively studied in the literature (Dann et al. (2023); Sutton and Barto (2018); Ng et al. (1999); Singh et al. (2010)). The idea here is to use domain knowledge and design a different, dense reward whose optimal policy is the same as that of the original reward. However, the learning algorithm deployed finds it much easier to learn the optimal policy via the shaped reward. While the existing works (eg Ng et al. (1999)) show that reward shaping methods can provably help learning in the single agent setting, their results apply to the multi-agent problem of RMABs, where the Lagrangian relaxation decouples the value function across arms and allows us to focus on separate single-agent MDPs. In this work, we eschew the careful human design of the reward function, and allow an LLM to assume this role. Our thesis is that the world knowledge the LLM obtains during its training, along with reflection, equips it with the domain knowledge essential to design efficient reward functions. Automated design of such reward functions have been explored in the literature: Mataric (1994); Sorg et al. (2010); Zheng et al. (2020). Here, a meta-network is trained to learn the best rewards along with the policy network which is trained to learn the optimal policy for a given reward. In our work, reflection serves a similar purpose as training of the meta-network but with the number of steps being dramatically smaller and the output being in the form of code, making it interpretable.
Algorithm 1 DLM Algorithm
1:Input: Task $\ell$, feature $\mathbf{z}$, and code $\zeta$ context
2:Hyperparameters: Loop iterations $I$, proposal batch $K$, sim. epochs $n_{\text{e}}$, sim. steps $n_{\text{s}}$, num. arms $N$
3:for iteration $=1$ to $I$ do
4:     ## LLM candidate reward proposal
5:     $R_{1:K}\sim\text{LLM}(\ell,\mathbf{z},\zeta)$
6:     Init. reflection string $\mathcal{Y}$
7:     ## Multi-agent simulations
8:     for reward $i=1$ to $K$ do
9:         Init. policy and critic nets. $\theta_{n},\phi_{n}\quad\forall n\in[N]$
10:         Init. $\lambda$-net. $\Lambda$, buffer $\mathcal{D}$, states $\mathbf{s}=\mathbf{s}_{0}$, features $\mathbf{z}$
11:         Clear reflection string $\mathcal{Y}=[]$
12:         for epoch $=1$ to $n_{\text{e}}$ do
13:              $\lambda=\Lambda(\mathbf{s})$
14:              for timestep $t=1$ to $n_{\text{s}}$ do
15:                  $a_{n}\sim\theta_{n}(s_{n},\lambda)\quad\forall n\in[N]$
16:                  $\mathbf{s^{\prime}},\mathbf{r}=\text{Simulate}(\mathbf{s},\mathbf{a},R_{i},\mathbf{z})$
17:                  $\mathcal{D}\leftarrow\mathcal{D}\cup\{(\mathbf{s},\mathbf{a},\mathbf{r},\mathbf{s^{\prime}},\lambda)\}$
18:                  $\mathbf{s}\leftarrow\mathbf{s^{\prime}}$
19:              Update each $(\theta_{n},\phi_{n})$ with PPO using $\mathcal{D}$
20:              Update $\Lambda$ with $\mathcal{D}$
21:         ## Outcome comparison
[…]
23:     ## Top candidate selection
[…]
25:     Update code context $\zeta\leftarrow\zeta\cup\{R_{\text{DLM}}\}$
26:Output: $R_{\text{DLM}}$
4.3 Provided DLM Context
In the reward candidate generation phase, we prompt the LLM with three key contextual components (shown Alg. 1, line 5). First, we provide a human-written language command that elucidates desired policy outcome. In practice, this is the only human input required for the proposed DLM technique.
Second, we provide arm feature category information, denoted $\mathbf{z}$. In the public health setting, features may include demographic information, socioeconomic data, or othe descriptors relevant to policy outcomes. We propose that LLMs may reasonably interpret the contextual relevance of respective features from a given language prompt, and adeptly propose reward functions using critical feature information to converge efficiently on desired policy outcomes. That is, given a language prompt $\ell$ we consider that LLMs may be used to extract the directly relevant features $\mathbf{z}_{dir}\subseteq\mathbf{z}$ explicitly referenced in $\ell$ to design a reward function $R$.
Third, we provide context regarding the RMAB problem formulation and corresponding code implementation. Specifically, we include information describing the state variable as a per-arm binary value, as well as syntax for accessing feature information from the codebase. From this context, we evaluate the ability for the LLM to propose reward functions in the form of code, as previously proposed by Ma et al. (2024); Li et al. (2023).
4.4 DRL Simulation
We evaluate each LLM-proposed reward function $R_{1:K}$ by training policy networks $\theta^{i}_{n}$ under each proposed reward $R_{i}$ for each arm in simulation $\forall n\in[N]$ (shown Alg. 1, lines 8:20). We consider a simulation space which defines the number of arms $N$, each with fixed, but hidden, transition dynamics. Following the procedure defined in Alg. 1, for a given reward $R_{i}$, we first sample an action-charge $\lambda$, which is used to decouple the learning of arm policies (Eq. 3, Alg. 1 line 13). Then, we simulate $n_{s}$ timesteps, sampling trajectories for each arm under the action-charge $\lambda$ (Alg. 1 lines 14:16). These trajectories are used in a buffer $\mathcal{D}$ (Alg. 1 line 17) to eventually update the policy and critic networks for each arm $(\theta_{n},\phi_{n})$, as proposed in Killian et al. (2022) (Alg. 1 lines 19:20). We compute advantage estimates for the actor from estimated Q-values.
Figure 2: Main results. We report interquartile mean Agarwal et al. (2021) and standard error. Task colors denote difficulty. Green tasks consider only one feature group (i.e. age, income, etc). Yellow tasks tasks consider more than one feature group, or single features not explicitly defined (e.g. Middle Income). Red tasks are highly ambiguous and require completely inferring the correct features (e.g. Early and Late Workers has no direct feature, requiring inference from preferred call times). Our method is able to match base reward performance in a completely automated fashion. We evaluate for a total $10$ seeds, each seed evaluating for $50$ separate trials and $10$ steps per trial, totaling $500$ samples of $10$ steps each per task. For each RMAB, we have a total of $48$ arms with a budget of $16$.
4.5 Reflection Stage
We propose the reflection comparison stage to enable LLM self-reflection and iteration on proposed rewards. At a high level, the outcome comparison produces a string showing, under each proposed reward $R_{i}$, the final simulated distribution of reward across key feature groups. Distributions, for instance, may be shown over age, gender, income, or other pre-defined feature characteristic groupings. This comparison is used in place of a numerical, scalar feedback function. Prior works assume access to some base fitness functions for self-reflection ($F$ as defined above) Singh et al. (2009). In the resource-constrained public health domain, we do not assume we have access to this ground truth function and extend the reward input from a scalar to a distribution.
Thus, rather than providing feedback through the base fitness function, we instead show the distribution of states achieved across various feature groups $g\in\mathbf{G}$; for example, over age or income brackets. Intuitively, because we focus on resource allocation in the public health domain, we propose that simulated state distributions provide sufficient signal for LLM to analyze the efficacy of proposed rewards, and select top candidate examples for future in-context learning.
We provide an overview of the policy outcome comparison procedure in Alg. 2. Given trained policy and critic networks $\theta_{n},\phi_{n}$, final lambda value $\lambda$, and feature matrix $\mathbf{Z}\in\mathbb{R}^{N\times m}$, we first simulate over $n_{\text{s}}$ evaluation timesteps (Alg. 2 lines 7:11). Note that we store only accumulated agent states $\mathbf{S}$; we use accumulated per-agent states to later reflect on achieved allocation outcomes. We then compute the percentage of reward attributed to arms within each sub-group $g\in\mathbf{G}$ (Alg. 2 lines 13:16). Finally, we report final reward distributions in an output string. We then query an LLM to select the best candidate example given the original prompt $\ell$ and the outcome distribution string $\mathcal{Y}$ (Alg. 1 lines 22:24). This selected example is then added as context in the next iteration of reward generation (Alg. 1 line 25).
This approach has two key strengths. First, we enable greater transparency of our approach, by using reward distributions to evaluate the effectiveness of a policy under a specified reward $\pi^{R}$. By providing only simulated outcome distributions from simulation to the reflection LLM, we ensure that self-reflection focuses only on alignment between the stated language goal and the resource allocation outcomes of our public health settings, without assuming any access to ground truth reward. Second, we allow for greater flexibility; by investigating only the outcome distributions of the proposed rewards, rather than original policies themselves, we enable a model-agnostic approach to self-iteration in reward design.
5 Experimental Evaluation
We evaluate DLM in a simulated setting representative of real-world resource allocation for maternal healthcare. We collaborate with ARMMAN, a non-profit based in India that enrolls pregnant women into a voice messaging service sharing preventative care information. We use real-world data from a service quality improvement study of 44,000 mothers collected by ARMMAN in January 2022 to construct the environment’s transition probabilities. This program, specifically focused on low-income communities Killian et al. (2022), employs health workers to provide service calls to high-risk beneficiaries; however, the number of women enrolled in the program far outnumbers the number of health workers. As such, the aim is to identify beneficiaries to call to maximize engagement with the program, allowing for the most efficient allocation of available resources.
Algorithm 2 Outcome Analysis
1:Input: Trained policy, critic nets. $\theta_{n},\phi_{n}\forall n\in[N]$, action-charge $\lambda$, feature matrix $\mathbf{Z}\in\mathbb{R}^{N\times m}$
2:Hyperparameters: Sim. steps $n_{\text{s}}$
[…]
4:Init. evaluation totals $\mathbf{S}\leftarrow[0]\times N$
5:Init. percentage distribution $\mathbf{P}\in\mathbb{R}^{\left|\mathbf{G}\right|}$
6:Init. OutputString as an empty string
7:for timestep $t=1$ to $n_{\text{s}}$ do
8:     $a_{n}\sim\theta_{n}(s_{n},\lambda)\quad\forall n\in[N]$
9:     $\mathbf{s^{\prime}},\mathbf{r}=\text{Simulate}(\mathbf{s},\mathbf{a},R_{i},\mathbf{z})$
10:     $\mathbf{S}\leftarrow\mathbf{S}+\mathbf{s}$
11:     $\mathbf{s}\leftarrow\mathbf{s^{\prime}}$
12:## Step 2: Compute state outcome distributions
13:$\mathbf{G}\leftarrow$ Define feature groups within $\mathbf{Z}$
14:for each group $g\in\mathbf{G}$ do
15:     $\mathbf{P}(g)\leftarrow\frac{\sum_{i\in g}\mathbf{S}_{i}}{\sum\mathbf{S}}$
16:     Append to OutputString: “$g$: $\mathbf{P}(g)\%$ of reward; ”
17:return OutputString
5.1 Tasks, Baselines, and Metrics
We provide experimental results for 11 distinct resource allocation prompts, shown in Appendix, Table 1. We separate tasks into three broad difficulty levels: simple single-feature tasks, middle-complexity multiple-feature tasks, and challenging, highly ambiguous tasks requiring complete inference of relevant features (see Figure 2). For each prompt, we compute normalized reward over 10 seeds; we describe the normalization equation below. To evaluate our proposed Decision-Language Model (DLM), we compare against several baseline reward functions:
Random: This policy samples arms uniformly at random. Through normalization, we set random performance score to be 0, as our floor. This would be equivalent to a maximum entropy reward, or simply a constant non-negative reward for any action. No Action: This policy samples no arms at each timestep. In most cases this will yield a negative normalized score, although in certain sparse reward tasks this may be close to the random policy. Default: This policy is trained with the same underlying algorithm as DLM, but the reward function is set to prioritize all beneficiaries equally; thus, for each arm, reward=1 for a positive state and reward=0 for a negative state. Most prior work in RMABs assume this reward. Base: This baseline sets the reward equal to the ground truth fitness function that all other methods are evaluated against. We use this reward as the consistent standard against which we measure the performance of our proposed method and other baselines. To be clear, each method uses a different reward at train time, but uses Base reward at evaluation time. We normalize with Base so that normalized reward is 1 for each task. DLM(No Reflection): This version of our proposed method ablates the reflection and iterative refinement process. It allows us to assess the importance of the reflection loop in improving the quality of the LLM-generated reward functions. DLM(Reflection): This is the full version of our proposed method.
[…]
5.2 Training Details
We use the Gemini Pro LLM model Team et al. (2023) from Google to generate reward functions, and train our downstream policy with RL using PPO Schulman et al. (2017) and the generated reward function. The downstream policy is trained for $2$ epochs total to prevent overfitting. We do a total of $2$ rounds of reflection and generate $2$ candidate reward functions per round. As detailed in Figure 2, we evaluate for a total $10$ seeds, each seed evaluating for $50$ separate trials and $10$ steps per trial, totaling $5000$ total steps per task.
5.3 DLM Success and Failure Cases
We present the experimental evaluation of DLM in Figure 2. We find that the proposed DLM technique is a capable reward designer across a broad range of tasks in public health, demonstrating, to the best of our knowledge, first example of LLMs as capable policy designers for resource allocation tasks in public health. Note that we evaluate under a sample-limited regime; thus, results demonstrate the ability to converge under limited samples to an optimal policy for a given prompt. Given this setting, we make several observations on the results in Figure 2.
Achieving base reward performance: We find DLM is able to reach the achieved base reward performance across a broad range of tasks. In particular, we find that the reflection process is able to improve the performance of the LLM in 9 out of 11 tasks, nearly reaching base reward in many circumstances. We find, in many cases, that LLMs can indeed identify the relevant features to use in proposed reward functions. An example can be seen for the task of biasing toward either tail of the age distribution (Age Distribution Tail Emphasis), from which the LLM is able to generate the base reward zeroshot:
 state * (agent_feats.youngest_agegroup or agent_feats.oldest_agegroup)
Benefits of reflection: We observe that the proposed DLM reflection technique, which grounds LLM reward generation by incorporating feedback from RMAB simulations, is able to consistently align policies closer to desired human preferences. We note, however, that the degree of improvement from reflection may vary by task type. For instance, in cases with extremely simple prompts, for example single feature group preferences, reward reflection may not substantially enhance allocation outcomes over the zeroshot LLM-proposed rewards (i.e. Age Distribution Tail Emphasis above).
Additionally, in highly ambiguous cases, such as Technologically Challenged, we find that reflection may not have sufficient prompt guidance to meaningfully iterate on proposed rewards. Further, for cases with extreme ambiguity, requiring high degrees of reasoning, we found that the LLM response would frequently state: ”There was not enough information specified to solve the task”; in these cases, no zeroshot reward is proposed, obviating the benefits of reflection. One potential solution, we suggest, may be to add a human in the proposal loop, to further specify prompts and refine the reward function in highly ambiguous cases. We leave these potential strategies for addressing highly ambiguous tasks to future work.
Misguided LLM reward proposals: We note that in certain cases, proposed rewards may have extremely poor performance prior to reflection (i.e. task Older Bias). In these cases, we observe that, while a task may be semantically simple, the LLM may attempt to use multiple features improperly. For example, for Older Bias, the LLM proposed:
state * agent_feats.secondoldest_agegroup * agent_feats.oldest_agegroup
Above, the LLM incorrectly uses features in the reward function, as an arm cannot be in both the second-oldest and the oldest age group simultaneously. Thus, in cases where improperly designed rewards are proposed, reflection may be helpful (see Appendix E). For more examples of generated reward functions, please see Appendix F.
Middle-complexity tasks: We observe that the reflection procedure may be most consistently helpful for the “middle range” of tasks that involve a composition of feature groups within the reasoning capabilities of the LLM. This may be particularly helpful in cases with complex goals that cannot easily be translated into updated policies by public health practitioners. For example, in ARMMAN’s Kilkari program (a national level program in India, or operational across 19 states in India) Bashingwa et al. (2021), different regions may have slightly different policy requirements, which a state program manager can achieve directly through language interfacing with DLM.
5.4 DLM Reflection Example
We visualize for two prompts, two candidate reward distributions and the selected best index in Figure 3. We show 4 relevant feature categories of age, income, education, and phone owners for both prompts. We observe that the DLM selection is able to reasonably interpret commands, such as prioritizing older ages, and select the distribution that best achieves the given prompt. However, we note that for more ambiguous criteria, such as “technologically challenged”, DLM selection may be more opaque; while the base reward for this prompt considers only income and phone ownership, additional features such as education level may also be considered for selection. In these cases, we hypothesize that additional human feedback may be required to better disambiguate prompts for improved reflection. We provide the full reward reflection text for the below examples in Appendix E.
Figure 3: DLM reflection distribution and selection visualization. DLM uses outcome distributions from trained policy simulations to refine proposed rewards. Above, we show the selected best candidate reward for two prompts. DLM is able to reasonably interpret given prompts using distributions across feature groups; however, for more ambiguous prompts (“technologically challenged”), the selection criteria may become more opaque, potentially requiring additional human feedback.
5.5 Exceeding Base in Critical Low Income
We find that empirically for the “Critical Low Income” task the LLM is actually able to improve over the base reward. We hypothesize here that the DLM model may propose a reward that is of much larger scaling than the original reward. For a representative example, we show a sample reward generated by the LLM for the Critical Low Income prompt:
(100 * state * agent_feats.no_or_lowest_income) + (100 * state * agent_feats.second_lowest_income) + state
In comparison, this is the corresponding base reward:
state + if_(state) * 10 * if_(agent_feats.no_or_lowest_income or agent_feats.second_lowest_income or agent_feats.third_lowest_income),
As the results reported in Figure 2 are taken in the low finite-sample regime, where there is not much data to train from, sample efficiency becomes extremely important. In this case, because the reward function generated by the LLM gives an order of magnitude higher weighting to specific arms, it seems likely that the underlying policy gradient algorithm PPO gains a large amount of sample efficiency in our setting. It is important to note that the improvement that the LLM yields over the base reward will shrink as the number of samples taken during training goes to infinity, as by definition the base reward defines the optimal policy. However, the highly shaped reward given by the LLM in this setting helps PPO converge faster for this task. It is also important to note that the base reward may not be the most sample efficient reward; other more complex rewards relying on correlated features or with different reward scaling are likely more efficient. Thus, this opens the possibility of achieving potential superhuman performance on our task in the future.
6 Conclusion
We propose a Decision-Language Model for resource allocation tasks in public health. The proposed model can take language prompts describing complex, potentially ambiguous policy outcome goals as input, and generate downstream control policies specialized for such goals. The proposed LLM-based approach can accurately capture the referenced prompt features and design effective reward functions in the form of code. Critically, our algorithms enable self-iteration on proposed reward functions in a fully automated fashion. By grounding LLM reward proposals with outcome distributions from RMAB simulations, we demonstrate the ability for the proposed DLM model to achieve near-base performance across a variety of policy preference tasks.
Ethical Statement
Consent is obtained from every beneficiary enrolling in the ARMMAN’s mobile health program (see also Appendix A). For future deployment of the proposed algorithm, one could take extra steps to avoid unsafe policies, including: (1) examining per beneficiary policy outcomes and performing spot-checks on trajectories; (2) apply the proposed algorithm to a subgroup of the population; (3) provide additional context to LLM after observing the impact on different features. We discuss measures to address fairness in Appendix B.
Appendix
Appendix A Data usage, consent, and secondary analysis
The data is collected from a service quality improvement study of 44K mothers conducted by ARMMAN in January 2022, and consent is obtained from every beneficiary enrolling in the NGO’s mobile health program. The data collected through the program is owned by the NGO and only the NGO is allowed to share data. The data exchange and usage were regulated by clearly defined exchange protocols including anonymization, read-access only to researchers, restricted use of the data for research purposes only, and approval by ARMMAN’s ethics review committee. No personally identifiable information (PII) is available to authors. We use anonymized call listenership logs to calculate empirical transition probabilities in experiments. Our experiment falls into secondary analysis of the data shared by ARMMAN and does not involve the deployment of the proposed algorithm or any other baselines to the service call program.
Appendix B Addressing Fairness
When LLM used is not well-aligned or calibrated to human values, there may be issues of fairness or bias. Although this work is a secondary analysis does not involve deployment of propose algorithms, future works on field studies may imposing fairness constraints on the reward function generation and use automatic detection of unfairness Dwork et al. [2012] to incorporate fairness. For future works on deployment, to ensure safe and responsible deployment, it would be interesting to study how to detect given prompts that are designed similarly to a leading question to induce hidden biases.
Appendix C Deep RL for RMABs
The Lagrangian relaxation allows us to disentangle the learning of arm policies and an action charge $\lambda$.
To learn arm policies, it is common to use RL approaches such as tabular Q-learning, deep Q-learning, or policy gradient algorithmsNakhleh et al. [2021]. In this paper, we use PPO Schulman et al. [2017], a policy gradient algorithm with an actor-critic structure. Specifically, the Q-function can be updated as follows:
$Q^{t}_{n}(s_{n},a_{n})\leftarrow$ $(1-\alpha_{q})Q_{n}^{t-1}(s_{n},a_{n})$
$+\alpha_{q}\left(R(s_{n})-\lambda c_{a_{n}},\beta\max_{a}Q(s_{n}^{\prime},a)\right),$
where $\alpha_{q}$ is the learning rate.
For an actor-critic methods, after performing an update of the Q-function in the critic, one may compute the advantage estimate:
$A\left(\mathbf{s},\mathbf{a}\right)=Q\left(\mathbf{s},\mathbf{a}\right)-V\left(\mathbf{s}\right).$
In practice, the values $A(s,a),Q(s,a),V(s)$ will be evaluated based on the current policy $\pi$. Intuitively, the advantage estimate tells us how much better $\mathbf{a}$ is compared to the current policy $\pi$. The advantage estimate is then plugged into the policy gradient to update the actor.
To learn an action charge $\lambda$, we follow an updating rule Killian et al. [2022]:
$\Lambda_{t}\leftarrow$ $\Lambda_{t-1}-\alpha_{\Lambda}\left(\frac{B}{1-\beta}+\sum_{n=1}^{N}\sum_{t=0}^{H}\mathbb{E}\beta^{t}c_{n,t}\right),$
where $c_{n,t}$ is the cost of the action taken by the optimal policy on arm $n$ in round $t$, and $\alpha_{\Lambda}$ is the step size.
Appendix D Provided Tasks
We show provided tasks in Table 1. We note that we evaluate on a wide variety of tasks within the real of resource allocation for maternal health. In particular, we test across feature characteristics such as age, income, spoken languages (an important characteristic indicating ethnicity in the Indian subcontienent), education, and other relevant features that may be considered in real-world allocation tasks. We provide these tasks directly as context to the LLM, in addition to relevant feature characteristics and their corresponding descriptions. Per seed, we fix the randomness in the sampled set of 48 arms from ARMMAN’s dataset, the deep RL neural network initialization, batch sampling from the dataset, and the sampling of states for each arm during evaluation. However please note that we cannot control the randomness of the Gemini LLM generation as we only have access through the Google API interface.
Table 1: Full prompt list
Label Task Command Base Function No Action Base Human Default DLM (No Reflection) DLM (Reflection)
Afternoon Call + NGO Registered
Advantage those who prefer being called between 12:30pm-3:30pm and are registered at an NGO.
state*.1 + if_(state) * 10 * if_(feature[29-1] and feature[33-1])
-0.50 ± 0.01
1.00 ± 0.07
-0.21 ± 0.06
0.88 ± 0.05
0.44 ± 0.07
0.82 ± 0.06
Low Income + Younger Bias
Combine the distributions of ‘age’ and ‘education’ to heavily give precedence to low income impoverished youth.
state*.1 + if_(state) * 10 * if_(feature[8-1] and feature[17-1])
-0.47 ± 0.00
1.00 ± 0.02
0.54 ± 0.03
0.42 ± 0.15
0.56 ± 0.02
0.96 ± 0.02
Marathi Speakers + Middle Aged
Fixate on those Marathi-speakers with middle-aged folks.
state*.1 + if_(state) * 10 * if_(feature[14-1] and (feature[10-1] or feature[11-1]))
-0.54 ± 0.01
1.00 ± 0.06
0.38 ± 0.06
0.68 ± 0.02
0.77 ± 0.05
0.86 ± 0.04
Older Bias
Focus on the most elderly by age distribution.
state*.1 + if_(state) * 10 * feature[12-1]
-0.70 ± 0.05
1.00 ± 0.18
0.48 ± 0.16
0.78 ± 0.01
-0.60 ± 0.16
0.71 ± 0.17
Age Distribution Tail Emphasis
Focus only on both the young and elderly.
state * if_(feature[12-1] or feature[8-1])
-1.19 ± 0.04
1.00 ± 0.15
0.73 ± 0.14
0.82 ± 0.03
1.13 ± 0.11
1.09 ± 0.17
Early and Late Workers
Give more attention for beneficiaries who likely work early in the morning and late at night.
state*.1 + if_(state) * 10 * if_(feature[27-1] or feature[29-1])
-0.60 ± 0.00
1.00 ± 0.03
0.84 ± 0.03
0.75 ± 0.03
0.63 ± 0.03
0.97 ± 0.04
Techincally Challenged
Infer technical challenges in reaching the phone that could indicate ‘at-risk’ beneficiaries and give preference.
state*.1 + if_(state) * 10 * if_((feature[25-1] or feature[26-1])) + if_(state) * 10 * 0.2 * if_((feature[37-1] or feature[38-1]))
-0.82 ± 0.03
1.00 ± 0.14
0.55 ± 0.11
0.66 ± 0.02
1.23 ± 0.15
1.00 ± 0.19
Middle Income
Prefer the income_bracket bounds for the middle 40% of the population.
state*.1 + if_(state) * 10 * if_(feature[39-1] or feature[40-1] or feature[41-1])
-0.65 ± 0.00
1.00 ± 0.03
0.50 ± 0.03
0.95 ± 0.02
0.37 ± 0.02
1.01 ± 0.03
Hindi Speakers
Select solely those who speak Hindi to analyze how this would affect the calls made to our beneficiaries.
state*.1 * if_(feature[13-1])
-0.62 ± 0.01
1.00 ± 0.04
0.24 ± 0.04
0.87 ± 0.01
0.67 ± 0.06
0.81 ± 0.07
Low Income
The lowest income_bracket distribution tail is very important, focus heavily on the top 12.5%.
state + if_(state) * 10 * feature[37-1]
-0.47 ± 0.00
1.00 ± 0.04
0.11 ± 0.05
0.72 ± 0.03
0.74 ± 0.04
0.96 ± 0.04
Critical Low Income
Weigth the lowest income_bracket groups, the lowest earners in the population.
state + if_(state) * 10 * if_(feature[36-1] or feature[37-1] or feature[38-1])
-0.66 ± 0.00
1.00 ± 0.03
0.33 ± 0.04
0.71 ± 0.03
0.32 ± 0.02
1.46 ± 0.03
Appendix E Sample Reward Reflection Outputs
Here we provide the full output of two sample reward reflection procedures, in addition to the corresponding selected LLM candidate reward.
E.1 Task: Older Age Bias
My goal was to create a Python reward function for RL in resource allocation,
with the objective of: Focus on the most elderly by age distribution. I tried
several reward functions for this task. Below, I have the given reward function,
and the corresponding distribution of reward achieved across 44 agent features.
A description of the features is as follows:
Index Name DataType
[sensitive features hidden]
 7. Ages 10-20 - Binary
 8. Ages 21-30 - Binary
 9. Ages 31-40 - Binary
 10. Ages 41-50 - Binary
 11. Ages 51-60 - Binary
 12. Speaks Hindi - Binary
 13. Speaks Marathi - Binary
 14. Speaks Gujurati - Binary
 15. Speaks Kannada - Binary
 16. Education level 1/7 -- illiterate - Binary
 17. Education level 2/7 -- 1-5th Grade Completed - Binary
 18. Education level 3/7 -- 6-9th Grade Completed - Binary
 19. Education level 4/7 -- 10th Grade Passed - Binary
 20. Education level 5/7 -- 12th Grade Passed - Binary
 21. Education level 6/7 -- Graduate - Binary
 22. Education level 7/7 -- Post graduate - Binary
 23. Phone owner 0 (e.g., woman) - Binary
 24. Phone owner 1 (e.g., husband) - Binary
 25. Phone owner 2 (e.g., family) - Binary
 26. To be called from 8:30am-10:30am - Binary
 27. To be called from 10:30am-12:30pm - Binary
 28. To be called from 12:30pm-3:30pm - Binary
 29. To be called from 3:30pm-5:30pm - Binary
 30. To be called from 5:30pm-7:30pm - Binary
 31. To be called from 7:30pm-9:30pm - Binary
 32. NGO - Binary
 33. ARMMAN - Binary
 34. PHC - Binary
 35. Income bracket -1 (no income) - Binary
 36. Income bracket 1 (e.g., 0-5000) - Binary
 37. Income bracket 2 (e.g., 5001-10000) - Binary
 38. Income bracket 3 (e.g., 10001-15000) - Binary
 39. Income bracket 4 (e.g., 15001-20000) - Binary
 40. Income bracket 5 (e.g., 20001-25000) - Binary
 41. Income bracket 6 (e.g., 25001-30000) - Binary
 42. Income bracket 7 (e.g., 30000-999999) - Binary
Below are the reward functions I used and their corresponding reward
distributions:
Index 0:
Reward Function:
state*agent_feats[10]+state+(state*agent_feats[10]==0)*agent_feats[11]
Reflection:
[sensitive features hidden]
Category: Ages
Ages 10-20: 1.30%
Ages 21-30: 13.46%
Ages 31-40: 10.46%
Ages 41-50: 2.57%
Ages 51-60: 72.23%
Category: Income
Income bracket -1 (no income): 0.00%
Income bracket 1 (e.g., 0-5000): 7.38%
Income bracket 2 (e.g., 5001-10000): 10.04%
Income bracket 3 (e.g., 10001-15000): 79.31%
Income bracket 4 (e.g., 15001-20000): 1.32%
Income bracket 5 (e.g., 20001-25000): 1.29%
Income bracket 6 (e.g., 25001-30000): 0.00%
Income bracket 7 (e.g., 30000-999999): 0.67%
Category: Calling Times
8:30am-10:30am: 79.52%
10:30am-12:30pm: 2.65%
12:30pm-3:30pm: 7.85%
3:30pm-5:30pm: 3.33%
5:30pm-7:30pm: 2.68%
7:30pm-9:30pm: 3.98%
Category: Education Levels
Illiterate: 2.80%
1-5th Grade Completed: 2.56%
6-9th Grade Completed: 5.37%
10th Grade Passed: 9.15%
12th Grade Passed: 3.96%
Graduate: 0.67%
Post graduate: 75.49%
Category: Languages Spoken
Speaks Hindi: 90.92%
Speaks Marathi: 9.08%
Speaks Gujurati: 0.00%
Speaks Kannada: 0.00%
Category: Phone Owners
Phone owner - Wom.: 98.59%
Phone owner - Hus.: 1.41%
Phone owner - Fam.: 0.00%
Category: Organizations
NGO: 90.70%
ARMMAN: 9.30%
PHC: 0.00%’
Index 1:
Reward Function: state * agent_feats[10] * agent_feats[11]
Reflection:
[some sensitive features hidden]
Category: Ages
Ages 10-20: 4.45%
Ages 21-30: 49.12%
Ages 31-40: 37.26%
Ages 41-50: 9.18%
Ages 51-60: 0.00%
Category: Income
Income bracket -1 (no income): 0.00%
Income bracket 1 (e.g., 0-5000): 23.70%
Income bracket 2 (e.g., 5001-10000): 42.50%
Income bracket 3 (e.g., 10001-15000): 24.82%
Income bracket 4 (e.g., 15001-20000): 4.48%
Income bracket 5 (e.g., 20001-25000): 4.50%
Income bracket 6 (e.g., 25001-30000): 0.00%
Income bracket 7 (e.g., 30000-999999): 0.00%
Category: Calling Times
8:30am-10:30am: 25.53%
10:30am-12:30pm: 11.45%
12:30pm-3:30pm: 26.00%
3:30pm-5:30pm: 13.50%
5:30pm-7:30pm: 11.58%
7:30pm-9:30pm: 11.94%
Category: Education Levels
Illiterate: 10.12%
1-5th Grade Completed: 8.83%
6-9th Grade Completed: 26.09%
10th Grade Passed: 29.74%
12th Grade Passed: 11.35%
Graduate: 2.27%
Post graduate: 11.60%
Category: Languages Spoken
Speaks Hindi: 66.41%
Speaks Marathi: 33.59%
Speaks Gujurati: 0.00%
Speaks Kannada: 0.00%
Category: Phone Owners
Phone owner - Wom.: 95.01%
Phone owner - Hus.: 4.99%
Phone owner - Fam.: 0.00%
Category: Organizations
NGO: 71.38%
ARMMAN: 28.62%
PHC: 0.00%’
Based on the above reward distributions and the given goal: Focus on the most
elderly by age distribution., please identify the index of the most effective
reward function. Provide your answer EXACTLY IN the following format: The best
reward function is at index: [INDEX]’.
The best reward function is at index: 0
E.2 Task: Technologically Challenged
My goal was to create a Python reward function for RL in resource allocation,
with the objective of: Infer technical challenges in reaching the phone that
could indicate at-risk beneficiaries and give preference. I tried several
reward functions for this task. Below, I have the given reward function, and the
corresponding distribution of reward achieved across 44 agent features. A
description of the features is as follows:
Index Name DataType
[sensitive features hidden]
 7. Ages 10-20 - Binary
 8. Ages 21-30 - Binary
 9. Ages 31-40 - Binary
 10. Ages 41-50 - Binary
 11. Ages 51-60 - Binary
 12. Speaks Hindi - Binary
 13. Speaks Marathi - Binary
 14. Speaks Gujurati - Binary
 15. Speaks Kannada - Binary
 16. Education level 1/7 -- illiterate - Binary
 17. Education level 2/7 -- 1-5th Grade Completed - Binary
 18. Education level 3/7 -- 6-9th Grade Completed - Binary
 19. Education level 4/7 -- 10th Grade Passed - Binary
 20. Education level 5/7 -- 12th Grade Passed - Binary
 21. Education level 6/7 -- Graduate - Binary
 22. Education level 7/7 -- Post graduate - Binary
 23. Phone owner 0 (e.g., woman) - Binary
 24. Phone owner 1 (e.g., husband) - Binary
 25. Phone owner 2 (e.g., family) - Binary
 26. To be called from 8:30am-10:30am - Binary
 27. To be called from 10:30am-12:30pm - Binary
 28. To be called from 12:30pm-3:30pm - Binary
 29. To be called from 3:30pm-5:30pm - Binary
 30. To be called from 5:30pm-7:30pm - Binary
 31. To be called from 7:30pm-9:30pm - Binary
 32. NGO - Binary
 33. ARMMAN - Binary
 34. PHC - Binary
 35. Income bracket -1 (no income) - Binary
 36. Income bracket 1 (e.g., 0-5000) - Binary
 37. Income bracket 2 (e.g., 5001-10000) - Binary
 38. Income bracket 3 (e.g., 10001-15000) - Binary
 39. Income bracket 4 (e.g., 15001-20000) - Binary
 40. Income bracket 5 (e.g., 20001-25000) - Binary
 41. Income bracket 6 (e.g., 25001-30000) - Binary
 42. Income bracket 7 (e.g., 30000-999999) - Binary
Below are the reward functions I used and their corresponding reward
distributions:
Index 0:
Reward Function:  2*state*(agent_feats[0]<36) + 3*state*(agent_feats[4]==1 or
agent_feats[4]==2) + 4*state*(agent_feats[5]<2)
Reflection:
  [sensitive features hidden]
Category: Ages
Ages 10-20: 1.97%
Ages 21-30: 8.36%
Ages 31-40: 89.25%
Ages 41-50: 0.42%
Ages 51-60: 0.00%
Category: Income
Income bracket -1 (no income): 0.00%
Income bracket 1 (e.g., 0-5000): 1.56%
Income bracket 2 (e.g., 5001-10000): 48.97%
Income bracket 3 (e.g., 10001-15000): 47.48%
Income bracket 4 (e.g., 15001-20000): 1.17%
Income bracket 5 (e.g., 20001-25000): 0.82%
Income bracket 6 (e.g., 25001-30000): 0.00%
Income bracket 7 (e.g., 30000-999999): 0.00%
Category: Calling Times
8:30am-10:30am: 2.75%
10:30am-12:30pm: 2.44%
12:30pm-3:30pm: 45.75%
3:30pm-5:30pm: 1.57%
5:30pm-7:30pm: 45.09%
7:30pm-9:30pm: 2.39%
Category: Education Levels
Illiterate: 0.00%
1-5th Grade Completed: 2.79%
6-9th Grade Completed: 48.13%
10th Grade Passed: 2.42%
12th Grade Passed: 3.70%
Graduate: 0.40%
Post graduate: 42.56%
Category: Languages Spoken
Speaks Hindi: 10.79%
Speaks Marathi: 89.21%
Speaks Gujurati: 0.00%
Speaks Kannada: 0.00%
Category: Phone Owners
Phone owner - Wom.: 16.37%
Phone owner - Hus.: 83.63%
Phone owner - Fam.: 0.00%
Category: Organizations
NGO: 95.53%
ARMMAN: 4.47%
PHC: 0.00%’
Index 1:
Reward Function:  -3.*(agent_feats[30] or agent_feats[31] or agent_feats[32] or
agent_feats[33] or agent_feats[34] or agent_feats[35] or agent_feats[36] or
agent_feats[37] or agent_feats[38] or agent_feats[39]) + state*1.
Reflection:
 [sensitive features hidden]
Category: Ages
Ages 10-20: 0.91%
Ages 21-30: 50.16%
Ages 31-40: 48.70%
Ages 41-50: 0.23%
Ages 51-60: 0.00%
Category: Income
Income bracket -1 (no income): 0.00%
Income bracket 1 (e.g., 0-5000): 0.66%
Income bracket 2 (e.g., 5001-10000): 72.25%
Income bracket 3 (e.g., 10001-15000): 25.95%
Income bracket 4 (e.g., 15001-20000): 0.67%
Income bracket 5 (e.g., 20001-25000): 0.47%
Income bracket 6 (e.g., 25001-30000): 0.00%
Income bracket 7 (e.g., 30000-999999): 0.00%
Category: Calling Times
8:30am-10:30am: 1.10%
10:30am-12:30pm: 1.38%
12:30pm-3:30pm: 47.13%
3:30pm-5:30pm: 0.89%
5:30pm-7:30pm: 24.65%
7:30pm-9:30pm: 24.85%
Category: Education Levels
Illiterate: 0.00%
1-5th Grade Completed: 1.36%
6-9th Grade Completed: 48.56%
10th Grade Passed: 1.14%
12th Grade Passed: 25.36%
Graduate: 0.22%
Post graduate: 23.37%
Category: Languages Spoken
Speaks Hindi: 27.75%
Speaks Marathi: 72.25%
Speaks Gujurati: 0.00%
Speaks Kannada: 0.00%
Category: Phone Owners
Phone owner - Wom.: 8.81%
Phone owner - Hus.: 69.26%
Phone owner - Fam.: 21.93%
Category: Organizations
NGO: 97.52%
ARMMAN: 2.48%
PHC: 0.00%’
Based on the above reward distributions and the given goal: Infer technical
challenges in reaching the phone that could indicate at-risk beneficiaries and
give preference., please identify the index of the most effective reward
function. Provide your answer EXACTLY IN the following format: The best reward
function is at index: [INDEX]’.
The best reward function is at index: 1
Appendix F Generated Reward Functions
Here we give all generated reward functions seen in Table 1 from the LLM. We first give the task description, which is fed to the LLM, the corresponding ground truth base reward, and then the generated LLM rewards in order of seed. Note that we have processed the original raw indices into feature labels for interpretability.
F.1 Focus on the most elderly by age distribution.
Ground truth base: lambda agent_feats, state: state*.1 + if_(state) * 10.0 * agent_feats.age_51_to_60
0: state * (agent_feats.age_51_to_60 or (agent_feats.age_51_to_60 and agent_feats.days_to_first_call >= 10))
1: state * (agent_feats.age_51_to_60 or agent_feats.ages41_to_50)
2: state * agent_feats.age_51_to_60
3: 2 * state * (agent_feats.age_51_to_60 or (agent_feats.ages41_to_50 and agent_feats.to_be_called730pm_930pm))
4: state * agent_feats.ages41_to_50 or state * agent_feats.age_51_to_60
5: state * (agent_feats.age_51_to_60 or agent_feats.ages41_to_50) + 1
6: 2*state + 20 * (agent_feats.age_51_to_60 and state)
7: state * (agent_feats.ages41_to_50 or agent_feats.age_51_to_60)
8: state * (agent_feats.age_51_to_60 or (agent_feats.ages41_to_50 and agent_feats.age_51_to_60))
9: state * (agent_feats.age_51_to_60 or agent_feats.ages41_to_50)
F.2 The lowest income_bracket distribution tail is very important, focus heavily on the top 12.5%.
Ground truth base: lambda agent_feats, state: state + if_(state) * HEAVILY_MODIFIER * agent_feats.second_least_income
0: state * max(0, agent_feats.highest_income_bracket - agent_feats.no_or_least_income)
1: state * (agent_feats.third_highest_income_bracket or agent_feats.second_highest_income_bracket or agent_feats.highest_income_bracket)
2: state * max(0, agent_feats.null - 0.875) * (agent_feats.highest_income_bracket or agent_feats.second_highest_income_bracket or agent_feats.third_highest_income_bracket)
3: state * ((agent_feats.second_least_income and not (agent_feats.no_or_least_income)) or (agent_feats.third_lowest_income_bracket and not (agent_feats.second_least_income)) or (agent_feats.fourth_lowest_income_bracket and not (agent_feats.third_lowest_income_bracket)) or (agent_feats.middle_income_bracket and not (agent_feats.fourth_lowest_income_bracket)) or (agent_feats.third_highest_income_bracket and not (agent_feats.middle_income_bracket)) or (agent_feats.second_highest_income_bracket or agent_feats.highest_income_bracket))
4: state * (agent_feats.to_be_called730pm_930pm or agent_feats.no_or_least_income) or (agent_feats.second_least_income and agent_feats.third_lowest_income_bracket and agent_feats.fourth_lowest_income_bracket and agent_feats.middle_income_bracket and agent_feats.third_highest_income_bracket and agent_feats.second_highest_income_bracket and agent_feats.highest_income_bracket)
5: (agent_feats.highest_income_bracket and state * 8) - (agent_feats.no_or_least_income and state * 4)
6: state * (5*agent_feats.highest_income_bracket - (agent_feats.no_or_least_income and not(agent_feats.second_least_income or agent_feats.third_lowest_income_bracket or agent_feats.fourth_lowest_income_bracket or agent_feats.middle_income_bracket or agent_feats.third_highest_income_bracket or agent_feats.second_highest_income_bracket)))
7: 100000 * state * (agent_feats.third_highest_income_bracket or agent_feats.second_highest_income_bracket or agent_feats.highest_income_bracket)
8: state * 100 * agent_feats.highest_income_bracket**2
9: state * (agent_feats.second_highest_income_bracket and agent_feats.highest_income_bracket)
F.3 Select solely those who speak Hindi to analyze how this would affect the calls made to our beneficiaries.
Ground truth base: lambda agent_feats, state: state*.1 * if_(agent_feats.speaks_hindi)
0: state * agent_feats.speaks_hindi
1: state * (agent_feats.speaks_hindi and not (agent_feats.speaks_marathi or agent_feats.speaks_gujurati or agent_feats.speaks_kannada))
2: (2 * state * agent_feats.speaks_hindi) + (3 * state)
3: state * (agent_feats.speaks_hindi and (not agent_feats.speaks_marathi and not agent_feats.speaks_gujurati and not agent_feats.speaks_kannada))
4: state * agent_feats.speaks_hindi
5: state*(agent_feats.speaks_hindi)
6: state * agent_feats.speaks_hindi
7: 1.2 * state * agent_feats.speaks_hindi
8: state * (agent_feats.speaks_hindi and (agent_feats.speaks_hindi or (not agent_feats.speaks_hindi and not agent_feats.speaks_marathi and not agent_feats.speaks_gujurati and not agent_feats.speaks_kannada)))
9: state * 2 * agent_feats.speaks_hindi
F.4 Focus only on both the young and elderly.
Ground truth base: lambda agent_feats, state: state * if_(agent_feats.ages_51_to_60 or agent_feats.ages_10_to_20)
0: state * (agent_feats.ages10_to_20 or agent_feats.age_51_to_60)
1: state * (agent_feats.ages10_to_20 and agent_feats.age_51_to_60)
2: state * (agent_feats.ages10_to_20 or agent_feats.age_51_to_60)
3: state * (agent_feats.ages10_to_20 or (agent_feats.ages41_to_50 and agent_feats.age_51_to_60))
4: state * (agent_feats.ages41_to_50 or agent_feats.ages10_to_20)
5: state * ((agent_feats.age_51_to_60 and not agent_feats.ages10_to_20) or (agent_feats.ages10_to_20 and not agent_feats.age_51_to_60))
6: state * ((agent_feats.ages10_to_20 or agent_feats.ages41_to_50) and not(agent_feats.ages_21_to_30 or agent_feats.ages_31_to_40))
7: state*(agent_feats.ages10_to_20 and agent_feats.age_51_to_60)
8: state * ((agent_feats.ages10_to_20 and agent_feats.ages41_to_50) or (agent_feats.age_51_to_60 and agent_feats.ages41_to_50))
9: state * (agent_feats.ages10_to_20 or agent_feats.age_51_to_60)
F.5 Weigth the lowest income_bracket groups, the lowest earners in the population.
Ground truth base: lambda agent_feats, state: state + if_(state) * 10.0 * if_(agent_feats.least_income_bracket or agent_feats.second_least_income or agent_feats.third_lowest_income_bracket)
0: state * (1 - agent_feats.no_or_least_income - agent_feats.second_least_income)
1: state * (1-agent_feats.no_or_least_income) * (1-agent_feats.second_least_income) * (1-agent_feats.third_lowest_income_bracket) * (1-agent_feats.fourth_lowest_income_bracket) * (1-agent_feats.middle_income_bracket) * (1-agent_feats.third_highest_income_bracket) * (1-agent_feats.second_highest_income_bracket)
2: state * (not agent_feats.no_or_least_income or not agent_feats.second_least_income)
3: state * (agent_feats.no_or_least_income or agent_feats.second_least_income or agent_feats.third_lowest_income_bracket or agent_feats.fourth_lowest_income_bracket or agent_feats.middle_income_bracket or agent_feats.third_highest_income_bracket or agent_feats.second_highest_income_bracket or agent_feats.highest_income_bracket)
4: 100 * state * agent_feats.no_or_least_income
5: 25 + state * agent_feats.no_or_least_income + (50 * state * agent_feats.second_least_income)
6: state * (agent_feats.no_or_least_income or agent_feats.second_least_income or agent_feats.third_lowest_income_bracket or agent_feats.fourth_lowest_income_bracket or agent_feats.middle_income_bracket or agent_feats.third_highest_income_bracket)
7: (state * agent_feats.no_or_least_income) or (state * agent_feats.second_least_income)
8: state * (5 - (agent_feats.no_or_least_income or agent_feats.second_least_income))
9: max(1.0, state*(1.1 if agent_feats.no_or_least_income or agent_feats.second_least_income else agent_feats.third_lowest_income_bracket or agent_feats.fourth_lowest_income_bracket or agent_feats.middle_income_bracket or agent_feats.third_highest_income_bracket or agent_feats.second_highest_income_bracket or agent_feats.highest_income_bracket))
F.6 Combine the distributions of ’age’ and ’education’ to heavily give precedence to low income impoverished youth.
Ground truth base: lambda agent_feats, state: state*.1 + if_(state) * 10.0 * if_(agent_feats.ages_10_to_20 and agent_feats.education_level_1)
0: state * (agent_feats.no_or_least_income and ((agent_feats.ages10_to_20 or agent_feats.ages_21_to_30) and (agent_feats.education_level_1 or agent_feats.education_level_2)))
1: state * (agent_feats.no_or_least_income and agent_feats.education_level5 and agent_feats.ages10_to_20)
2: state * (agent_feats.ages41_to_50 and agent_feats.education_level_1 and agent_feats.no_or_least_income and agent_feats.days_to_first_call > 15) + state * (agent_feats.ages41_to_50 and agent_feats.education_level_2 and agent_feats.second_least_income and agent_feats.days_to_first_call > 15) + state * (agent_feats.ages41_to_50 and agent_feats.education_level_3 and agent_feats.third_lowest_income_bracket and agent_feats.days_to_first_call > 15)
3:(state and (agent_feats.no_or_least_income or (agent_feats.education_level_6 and agent_feats.ages10_to_20)))
4:state * (   (agent_feats.education_level_1 and agent_feats.no_or_least_income) or (agent_feats.education_level_2 and agent_feats.second_least_income) or (agent_feats.education_level_3 and agent_feats.third_lowest_income_bracket) or (agent_feats.education_level_4 and agent_feats.fourth_lowest_income_bracket)  )
5: state * (agent_feats.ages10_to_20 and (agent_feats.education_level_6 or agent_feats.education_level_7) and agent_feats.no_or_least_income)
6: 100 * state * (agent_feats.ages10_to_20 and (agent_feats.education_level_1 or agent_feats.education_level_3 or agent_feats.education_level5) and agent_feats.no_or_least_income)
7: 2*state*(agent_feats.no_or_least_income and (agent_feats.ages41_to_50 or agent_feats.age_51_to_60 or agent_feats.speaks_hindi) and (agent_feats.education_level_1 or agent_feats.education_level_2 or agent_feats.education_level_3))
8: state * (agent_feats.no_or_least_income and (agent_feats.education_level_1 or agent_feats.education_level_2 or agent_feats.education_level_3))
9: state * (agent_feats.no_or_least_income and ((agent_feats.ages10_to_20 or agent_feats.ages_21_to_30) and (agent_feats.education_level_1 or agent_feats.education_level_2 or agent_feats.education_level_3)))
F.7 Advantage those who prefer being called between 12:30pm-3:30pm and are registered at an NGO.
Ground truth base: lambda agent_feats, state: state*.1 + if_(state) * 10.0 * if_(agent_feats.called1030am_to_1230pm and agent_feats.ngo)
0: state * (agent_feats.ngo and agent_feats.called1230_to_330) * (agent_feats.ages10_to_20 or agent_feats.ages_21_to_30 or agent_feats.ages_31_to_40)
1: state * agent_feats.ngo * agent_feats.called1230_to_330 * agent_feats.age_51_to_60 * agent_feats.speaks_hindi
2: state * (agent_feats.called1230_to_330 and agent_feats.ngo)
3: state * (agent_feats.called1230_to_330 + (agent_feats.ngo and agent_feats.called1230_to_330))
4:(state and (agent_feats.called1230_to_330 or agent_feats.ngo)) + (2*state*agent_feats.called1230_to_330*agent_feats.ngo)
5:state * (agent_feats.called1230_to_330 and (agent_feats.ngo or not agent_feats.called1230_to_330))
6: state * (agent_feats.ngo and agent_feats.called1230_to_330) + state * (agent_feats.ngo and agent_feats.ages_21_to_30 and not agent_feats.ages_31_to_40 and not agent_feats.ages41_to_50)
7: state * (agent_feats.ngo and agent_feats.called1230_to_330 and (agent_feats.ages10_to_20 or agent_feats.ages_21_to_30 or agent_feats.ages_31_to_40 or agent_feats.ages41_to_50))
8: state * (agent_feats.ngo and agent_feats.called1230_to_330 and ((agent_feats.to_be_called830am_to_1030am or agent_feats.to_be_called1030am_1230pm) or (not agent_feats.to_be_called830am_to_1030am and not agent_feats.to_be_called1030am_1230pm)))
9:state * (agent_feats.ngo and agent_feats.called1230_to_330 and agent_feats.ages41_to_50 or (agent_feats.ngo and agent_feats.called1230_to_330))
F.8 Prefer the income_bracket bounds for the middle 40% of the population.
Ground truth base: lambda agent_feats, state: state*.1 + if_(state) * 10.0 * if_(agent_feats.fourth_lowest_income_bracket or agent_feats.middle_income_bracket or agent_feats.third_highest_incom_bracket)
0:state * (agent_feats.second_least_income or agent_feats.third_lowest_income_bracket or agent_feats.fourth_lowest_income_bracket or agent_feats.middle_income_bracket or agent_feats.third_highest_income_bracket)
1: state * ((agent_feats.second_least_income or agent_feats.third_lowest_income_bracket or agent_feats.fourth_lowest_income_bracket or agent_feats.middle_income_bracket) and not (agent_feats.no_or_least_income or agent_feats.third_highest_income_bracket or agent_feats.second_highest_income_bracket or agent_feats.highest_income_bracket))
2: state * (agent_feats.second_least_income or agent_feats.third_lowest_income_bracket or agent_feats.fourth_lowest_income_bracket or agent_feats.middle_income_bracket)
3: state * ((agent_feats.fourth_lowest_income_bracket or agent_feats.middle_income_bracket) and not (agent_feats.second_least_income or agent_feats.third_lowest_income_bracket or agent_feats.third_highest_income_bracket or agent_feats.second_highest_income_bracket or agent_feats.highest_income_bracket))
4: state * (agent_feats.second_least_income or agent_feats.third_lowest_income_bracket or agent_feats.fourth_lowest_income_bracket or agent_feats.middle_income_bracket)
5: state * (agent_feats.third_lowest_income_bracket and agent_feats.fourth_lowest_income_bracket and agent_feats.middle_income_bracket)
6: state * (agent_feats.third_lowest_income_bracket and agent_feats.fourth_lowest_income_bracket and agent_feats.middle_income_bracket and agent_feats.third_highest_income_bracket)
7:state * agent_feats.fourth_lowest_income_bracket
8: state * (agent_feats.second_least_income and agent_feats.third_lowest_income_bracket and agent_feats.fourth_lowest_income_bracket and agent_feats.middle_income_bracket and not(agent_feats.no_or_least_income or agent_feats.third_highest_income_bracket or agent_feats.second_highest_income_bracket or agent_feats.highest_income_bracket))
9: state * ((agent_feats.second_least_income and agent_feats.third_lowest_income_bracket and agent_feats.fourth_lowest_income_bracket and agent_feats.middle_income_bracket and not agent_feats.third_highest_income_bracket) or (agent_feats.second_least_income and agent_feats.third_lowest_income_bracket and agent_feats.fourth_lowest_income_bracket and agent_feats.middle_income_bracket and agent_feats.third_highest_income_bracket))
F.9 Fixate on those Marathi-speakers with middle-aged folks.
Ground truth base: lambda agent_feats, state: state*.1 + if_(state) * 10.0 * if_(agent_feats.speaks_marathi and (agent_feats.ages31_to_40 or agent_feats.ages41_to_50))
0: state * (agent_feats.speaks_marathi and (agent_feats.ages_21_to_30 or agent_feats.ages_31_to_40 or agent_feats.ages41_to_50))
1: state * (agent_feats.speaks_marathi and agent_feats.ages_31_to_40 and agent_feats.called530pm_to_730pm)
2: state * (agent_feats.speaks_marathi and agent_feats.ages_31_to_40)
3: 2 * state and (18 and agent_feats.speaks_marathi and agent_feats.ages_31_to_40 and agent_feats.phone_owner_husband)
4: state * (agent_feats.speaks_marathi and agent_feats.ages_31_to_40 and agent_feats.education_level_7)
5: state * (agent_feats.speaks_marathi and agent_feats.ages_31_to_40)
6: 2*state*(agent_feats.speaks_marathi and ((agent_feats.ages_31_to_40 or agent_feats.ages41_to_50) or (agent_feats.ages_21_to_30 and agent_feats.age_51_to_60)))
7: state * agent_feats.speaks_marathi * agent_feats.ages41_to_50
8: 1 + 5*state + 40*(state and agent_feats.speaks_marathi and agent_feats.ages_31_to_40) + 10*agent_feats.ages_21_to_30
9: state * (agent_feats.speaks_marathi and (agent_feats.ages_31_to_40 or agent_feats.ages41_to_50))
F.10 Give more attention for beneficiaries who likely work early in the morning and late at night.
Ground truth base: lambda agent_feats, state: state*.1+ if_(state) * 10.0 * if_(agent_feats.to_be_called830am_to1030am or agent_feats.called_1230pm_to_330pm)
0: state * (agent_feats.to_be_called830am_to_1030am or agent_feats.to_be_called730pm_930pm)
1: state * (agent_feats.to_be_called830am_to_1030am or agent_feats.to_be_called730pm_930pm)
2: state * (agent_feats.to_be_called830am_to_1030am or agent_feats.to_be_called730pm_930pm)
3: 10 * state * (agent_feats.to_be_called830am_to_1030am or agent_feats.to_be_called730pm_930pm)
4: state * (agent_feats.to_be_called830am_to_1030am and agent_feats.to_be_called730pm_930pm)
5: state * (agent_feats.to_be_called830am_to_1030am or agent_feats.to_be_called730pm_930pm)
6: state * ((agent_feats.called1230_to_330 or agent_feats.to_be_called330pm_to_530pm) and (agent_feats.to_be_called830am_to_1030am or agent_feats.to_be_called730pm_930pm))
7: state * (((agent_feats.to_be_called830am_to_1030am and agent_feats.to_be_called730pm_930pm) or (agent_feats.to_be_called1030am_1230pm and agent_feats.called1230_to_330)) or (agent_feats.to_be_called330pm_to_530pm and agent_feats.called530pm_to_730pm))
8: state * ((agent_feats.to_be_called830am_to_1030am or agent_feats.to_be_called730pm_930pm) or (agent_feats.to_be_called1030am_1230pm or agent_feats.called530pm_to_730pm) or (agent_feats.called1230_to_330 or agent_feats.to_be_called330pm_to_530pm))
9: state*(agent_feats.to_be_called830am_to_1030am or agent_feats.to_be_called730pm_930pm)
F.11 Infer technical challenges in reaching the phone that could indicate ’at-risk’ beneficiaries and give preference.
Ground truth base: lambda agent_feats, state: state*.1+ if_(state) * 10.0 * if_((agent_feats.phone_owner_husband or agent_feats.phone_owner_family))
0: state * (agent_feats.ages41_to_50 or agent_feats.ngo)
1: state * ((agent_feats.education_level_1 or agent_feats.education_level_2) and agent_feats.highest_income_bracket)
2: state * (agent_feats.speaks_hindi and agent_feats.highest_income_bracket and agent_feats.phone_owner_woman or agent_feats.speaks_gujurati and agent_feats.to_be_called830am_to_1030am and agent_feats.to_be_called730pm_930pm)
3: state * ((not agent_feats.ages10_to_20) and (agent_feats.speaks_hindi or agent_feats.speaks_marathi))
4: state * ((agent_feats.enrollment_gestational_age < 37) and (agent_feats.days_to_first_call > 15) or (agent_feats.to_be_called1030am_1230pm and agent_feats.speaks_marathi) or (agent_feats.enrollment_delivery_status and not agent_feats.ages_21_to_30))
5: state + 5 * (agent_feats.ages10_to_20 or agent_feats.education_level5 or agent_feats.to_be_called1030am_1230pm)
6: state * ((not agent_feats.phone_owner_woman) or agent_feats.days_to_first_call > 7)
7: state * ((agent_feats.education_level_2 or agent_feats.education_level_1) and (agent_feats.to_be_called1030am_1230pm or agent_feats.called1230_to_330))
8: state * (1.2*agent_feats.age_51_to_60 + (agent_feats.live_births_count == 0) + agent_feats.ngo + agent_feats.armman)
9: state*(agent_feats.enrollment_delivery_status and ((agent_feats.ages10_to_20 or agent_feats.ages_21_to_30 or agent_feats.ages_31_to_40 or agent_feats.ages41_to_50) or (agent_feats.phone_owner_woman and not agent_feats.to_be_called830am_to_1030am and not agent_feats.to_be_called1030am_1230pm)))